OpenStudy (anonymous):

The sales tax on a table saw is 6.96 What is the price of the table saw before tax if the sales tax rate is 5.8%

OpenStudy (anonymous):

6.96/5.8% will be the price of the table saw.

OpenStudy (anonymous):

@shanlee can you plz help me with a question?

OpenStudy (anonymous):

6.96/5.8% =6.96*100/5.8 =696/5.8 =120 120 is the price of the table saw.

OpenStudy (anonymous):

thankk youu

OpenStudy (anonymous):

You are welcome.